Bill and heather each improved their yards by planting rose bushes and geraniums. They bought their supplies from the same store. Bill spent $47 on 13 rose bushes and 7 geraniums. Heather spent $44 on 7 rose bushes and 10 geraniums. What is the cost for each plant?
Answer by Boreal(15235)

rose=x
geranium=y
13x+7y=47
7x+10y=44
multiply top by -10 and bottom by 7
-130x-70y=-470
49x+70y=308. Now add, and y s disappear
-81x=-162
x=$2 per rose bush
substitute into first and 26+7y=47
7y=21
y=$3 per geranium
Check in second
7(2)+10(3)=44
Works.